Five times the sum of a number and 3 is the same as 3 multiplied by 1 less than twice the number

Answer by algebrahouse.com(1659)

x = the number

5(x + 3) = 3(2x - 1) {five times sum of number and 3 equals 3 multiplied by 1 less than twice number}
5x + 15 = 6x - 3 {used distributive property}
x = 18 {subtracted 5x and added 3 to each side}

For more help from me, visit: www.algebrahouse.com